a. How many grams of aluminum are in 0.53 L of aluminum?


b.If there are 10.0 lbs of aluminum, how many Liters of aluminum are there?



1
Expert's answer
2020-09-17T07:15:49-0400

Formula which we will be using here

"Density=\frac{mass}{volume}"


Density of aluminum is 2.7 g/cm3

Volume given=0.53 L=530 cm3

Mass of aluminum=2.7×530=1431 g


1 lbs = 453.6 grams

10 lbs = 4536 grams

volume of aluminium = "\frac{4536}{2.7}=1680 cm^3"

converting this into litres = 1.68 L